if the force between the elctron inthe first bohr orbit and the nucleous in hydrogen atom is f , then the force between them when the electron is in second orbit is

Grade:12

1 Answers

Vikas TU
7 years ago
r1 = 0.0529 x 1^2/1
f  = k * e^2/r1 = >  ke^2/0.0529
 
for 2nd orbit,
r2 = 0.0529 x 2^2/1) = > 0.0529 x 4
F =ke^2/(0.0529 x 4)
Divide,
F/f we get, =>
F/f = (ke^2/(0.0529 x 4)) *(0.0529/ke^2)) = > ¼
f = 4F
or
F = f/4